OpenStudy (anonymous):

(9x ^ 4y ^ 8) ^ 2/(3x 3y 5) ^ 2

OpenStudy (tkhunny):

Have you considered applying the laws of exponents? \(\left(a^{m}\right)^{n} = a^{m\cdot n}\) \(\dfrac{b^{m}}{b^{n}} = b^{m-n}\) You'll be done in no time.

OpenStudy (anonymous):

our teacher isn't much help so understanding this is hard and I'm stuck like big time